DK4L:

An examination consists of three parts. In part A, a student must answer 2 of 3 questions. In part B, a student must answer 6 of 8 questions and in part C, a student must answer all questions. How many choices of questions does the student have?

DanGonz:

non nom non

DK4L:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @DanGonz non nom non \(\color{#0cbb34}{\text{End of Quote}}\) wat?

DanGonz:

the student has 4 choices

DanGonz:

ANYWAYS........

DK4L:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @DanGonz ANYWAYS........ \(\color{#0cbb34}{\text{End of Quote}}\) how is it 4

DanGonz:

hehe um iont know

DanGonz:

i guess

DK4L:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @DanGonz hehe um iont know \(\color{#0cbb34}{\text{End of Quote}}\) oh u want me to fail

DanGonz:

i never said that

DanGonz:

im tryin to be funny

DanGonz:

well it dont work bye

DK4L:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @DanGonz im tryin to be funny \(\color{#0cbb34}{\text{End of Quote}}\) ik

DanGonz:

:(

DK4L:

I just don't want to get it wrong

XioGonz:

Given: x2 - 3|x - 2| - 4x = - 6 Let Y = x - 2 which gives x = Y + 2 Substitute in the equation above: (Y + 2)2 - 3|Y| - 4(Y + 2) = - 6 Y2 - 3|Y| + 2 = 0 Notice: Y2 = |Y|2 Rewrite the equation as: |Y|2 - 3|Y| + 2 = 0 (|Y| - 2)(|Y| - 1) = 0 Solve for |Y|: |Y| = 2 , |Y| = 1 Solve for Y: Y = 2, -2 , 1 , -1 solve for x using x = Y + 2. So: x = 4 , 0 , 3 , 1